OpenStudy (anonymous):

How do we prove that lim of sigma 1/k is +oo

OpenStudy (lgbasallote):

\[\huge \sigma \frac 1k = +\infty\] ??

OpenStudy (shubhamsrg):

maybe you mean summation ( 1/ k) where k goes from 1 to infinity ?

OpenStudy (shubhamsrg):

on the LHS*

OpenStudy (anonymous):

the limite of the summation 1/k where k goes from 0 to n

OpenStudy (anonymous):

And n goes to infinity

OpenStudy (unklerhaukus):

\[\lim\limits_{n\rightarrow\infty}\quad \sum\limits_{k=0}^n\quad\frac{1}{k}\]

OpenStudy (unklerhaukus):

??

OpenStudy (anonymous):

Yes that's right. Please help

OpenStudy (unklerhaukus):

well just try the fist term , when k=0....

OpenStudy (anonymous):

Yeah but that's not possible. what if k goes from 1 to n, what do we do then?